Galaxy United, Inc.
2009 Income Statement
($ in millions)
  Net sales $8,550     
  Less: Cost of goods sold 7,180     
  Less: Depreciation

420     

  Earnings before interest and taxes 950     
  Less: Interest paid

78     

  Taxable Income 872     
  Less: Taxes

305     

  Net income

$ 567     

Galaxy United, Inc.
2008 and 2009 Balance Sheets
($ in millions)

2008

2009

2008

2009

  Cash $   120      $ 160         Accounts payable $1,120        $1,140    
  Accounts rec. 940      780         Long-term debt   940        1,204    
  Inventory

1,470     

1,510      

  Common stock $3,180      $2,980    
  Sub-total $2,530      $2,450         Retained earnings

  510     

776    

  Net fixed assets

3,220     

3,650      

  Total assets

$5,750     

$6,100      

  Total liab. & equity

$5,750    

$6,100    


What is the return on equity for 2009?

rev: 01_14_2016_QC_CS-37830

13 percent

16 percent

10 percent

18 percent

15 percent
Answer #1

Return On Equity (ROE) = Net income / Average shareholders Equity

Average shareholders Equity = (Beginning shareholders Equity + Ending shareholders Equity)/2

= (3180+510+2980+776)/2

= 7446/2

= 3723

Return On Equity (ROE) = 567/3723

= 15.23%

15 percent